# thanks: Brinson Prize Fellow

### v1

@@ Abstract
We present a measurement of the free-streaming length of dark matter (DM) and subhalo abundance around 28 quadruple image strong lenses using observations from JWST MIRI presented in Paper III of this series. We improve on previous inferences on DM properties from lensed quasars by simultaneously reconstructing extended lensed arcs with image positions and relative magnifications (flux ratios). Our forward modeling framework generates full populations of subhalos, line-of-sight halos, and globular clusters, uses an accurate model for subhalo tidal evolution, and accounts for free-streaming effects on halo abundance and concentration. Modeling lensed arcs leads to more-precise model-predicted flux ratios, breaking covariance between subhalo abundance and the free-streaming scale parameterized by the half-mode mass $m_{\rm{hm}}$. Assuming subhalo abundance predicted by the semi-analytic model galacticus ($N$-body simulations), we infer (Bayes factor of 10:1) $m_{\rm{hm}}<10^{7.4}\mathrm{M}_{\odot}$ ($m_{\rm{hm}}<10^{7.2}\mathrm{M}_{\odot}$), a 0.4 dex (0.3 dex) improvement relative to omitting lensed arcs. These bounds correspond to lower limits on thermal relic DM particle masses of $7.4$ and $8.4$ keV, respectively. Conversely, assuming DM is cold, we infer a projected mass in subhalos ($10^{6}<m/M_{\odot}<10^{10.7}$) of $1.6_{-1.1}^{+2.4}\times 10^{7}\ \mathrm{M}_{\odot}\ \rm{kpc^{-2}}$ at $95\%$ confidence. This is consistent with galacticus predictions ($0.6\times 10^{7}\mathrm{M}_{\odot}\ \rm{kpc^{-2}}$), but in tension with recent $N$-body simulations ($0.3\times 10^{7}\mathrm{M}_{\odot}\ \rm{kpc^{-2}}$). Our results are the strongest limits on WDM, and the most precise measurement of subhalo abundance around strong lenses. Further improvements will follow from the large sample of lenses to be discovered by Euclid, Rubin, and Roman.
@@ II.2.2 The flux ratio likelihood
We apply the ABC approach to one system in our sample, J0924, because the flux ratios deviate strongly from the predictions of a smooth lens model. As a result, the flux ratio likelihood heavily downweights the majority of lens models. To generate enough samples to fully explore the parameter space we use the ABC approximation of the exact flux ratio likelihood. With ABC, we can guarantee that we have enough samples to explore the parameter space. We accept the 5,000 samples corresponding to the smallest $S$ out of $\sim 6,000,000$ total lens models. This amounts to an implicit choice of $\epsilon$.
@@ II.3 Imaging data reconstruction
There is not a unique way to choose $\sigma_{\rm{noise}}$. We have experimented with defining this term in a way that connects directly to the uncertainties in the imaging data on a pixel level. For example, we can generate 10,000 background noise realizations (using $\sigma_{ii,(\rm{rms})}$ and $\sigma_{ii,\rm{(Poisson)}}$), render them on top of a best-fit lens model for each system, and set $\sigma_{ii,(\rm{noise})}$ equal to the standard deviations of the resulting distribution of log-likelihoods. However, strategies for defining $\sigma_{(\rm{noise})}$ in a way that connects to the background and Poison noise in the image tends to over-correct, in the sense that we discard more information from the imaging data that is necessary to make the problem computationally tractable. Instead, since the primary motivation for including $\sigma_{(\rm{noise})}$ is related to available CPU hours, $n_{\rm{cpu-hour}}$, we define $\sigma_{(\rm{noise})}$ in such a way that $\sigma_{(\rm{noise})}\rightarrow 0$ (meaning we do not down-weight constraints from lensed arcs) as $n_{\rm{cpu-hour}}\rightarrow\infty$. To compute $w_{\rm{img}}$, we generate $N\sim 200,000$ lens models per system, on average, and set $\sigma_{(\rm{noise})}$ such that $N/N_{\rm{eff}}=2000$. The threshold $N/N_{\rm{eff}}=2000$ represents a compromise between extracting more information from the lensed arcs, and having enough sampling of the parameter space to suppress shot-noise in the $w_{\rm{img}}$ distributions. We offer a complementary interpretation of the $\sigma_{(\rm{noise})}$ parameter: this parameter effectively bootstraps the uncertainties in the imaging data until the best 2000 lens models have comparable log-likelihoods.
@@ III Datasets
Table 1: Description of the dark matter hyper-parameters introduced in Section IV. The predictions for $\delta_{\rm{LOS}}$, $\alpha$ and $\Sigma_{\rm{sub}}$ remain unchanged in WDM, as these parameters determine the form of the subhalo mass function anchored at high masses. Log-uniform priors are in base 10.
@@ IV.1.2 Subhalos
Our strategy for modeling subhalo density profiles uses the tidal evolution model presented by Du++25. This model is designed specifically for the analysis of strong gravitational lens systems, and predicts the density profile of subhalos that appear in projection inside $3\theta_{\rm{E}}$ based on the subhalo properties at infall, given the infall subhalo mass function.
We begin by defining the differential subhalo infall mass function per unit area
@@
After using Equation 15 to generate the subhalo masses at infall, we use the tidal stripping framework presented by Du++25 to predict their density profiles. The calculation begins with assigning infall redshifts to each subhalo using the probability distribution $p\left(z_{\rm{infall}}|m,m_{\rm{host}}\right)$, where the infall redshift depends on $m$ and the host-halo mass $m_{\rm{host}}$. From $z_{\rm{infall}}$ and $m$ we evaluate the concentration of the subhalo at infall $c=c\left(m,z_{\rm{infall}}\right)$. Next, given the host halo concentration $c_{\rm{host}}=c_{\rm{host}}\left(m_{\rm{host}},z_{\rm{d}}\right)$, which we calculate using the DiemerJoyce19 concentration–mass relation, we sample the bound mass fraction, $f_{\rm{bound}}\equiv m_{\rm{bound}}/m$, from the probability distribution $p\left(f_{\rm{bound}}|c,z_{\rm{infall}},c_{\rm{host}}\right)$. Du++25 derived empirical representations of this distribution by evolving subhalos in a growing host potential with galacticus. As discussed by Du++25, a given set of subhalo infall time, concentration, and host halo concentration specifies a distribution of $f_{\rm{bound}}$ due to differences in subhalo orbits and mass loss histories. Given $f_{\rm{bound}}$, we use tidal tracks [e.g. Errani++21, Du++24] to obtain the subhalo density profile, the truncation radius $r_{\mathrm{t}}$ and a factor $f_{\rm{t}}$ that changes the normalization of the NFW profile $\rho_{\mathrm{s}}\rightarrow f_{\rm{t}}\rho_{\mathrm{s}}$. This model predicts that, at the redshift of the lens, most subhalos have lost over $90\%$ of the mass they had at infall. The model implemented in $\tt{pyHalo}$ shows excellent agreement with cosmological realizations of subhalo evolution performed with galacticus for $\Sigma_{\rm{sub}}\approx 0.1\ \rm{kpc^{-2}}$.
@@ IV.2 Extension to warm dark matter
The values of $k_{\rm{hm}}$, $\beta$, and $\gamma$ depend on the mass and formation mechanism of the dark matter particle. For example, a sterile neutrino of the same mass can be warm or cold, depending on the production mechanism [Kusenko:2009up, Abazajian:2019ejt]. To enable comparison with existing WDM constraints, for this initial analysis of our full JWST dataset we assume the same thermal relic WDM model considered in many previous works. For a thermal relic WDM particle [^15] For thermal relic WDM we have $\beta\sim 2$ and can take $\gamma=5$ in Equation 17 [Stucker++22]., we can relate the half-mode mass to the particle mass
@@
IV.3 Sampling distributions for dark matter hyper-parameters
We use a uniform sampling distribution on $\log_{10}m_{\rm{hm}}/\mathrm{M}_{\odot}$ in the range $\left[4,10\right]$, and quote our constraints on this parameter with the same Bayesian prior. While CDM predicts the existence of halos down to Earth-mass scales, and therefore $m_{\rm{hm}}=10^{4}\mathrm{M}_{\odot}$ is not formally consistent with CDM, for $m_{\rm{hm}}=10^{4}\mathrm{M}_{\odot}$ the abundance and density profiles of halos deviate by less than $10\%$ at $10^{6}\mathrm{M}_{\odot}$, the minimum halo mass rendered in our simulations. Therefore, we can identify regions of parameter space with $m_{\rm{hm}}<\mathcal{O}\left(10^{5}\mathrm{M}_{\odot}\right)$ as effectively indistinguishable from CDM.
We draw $\delta_{\rm{LOS}}$ (Equation 13) from a sampling distribution that is uniform in $\left[0.9,1.1\right]$, allowing for $10\%$ difference in the amplitude of the line-of-sight halo mass function. We draw $\alpha$ from a uniform sampling distribution in $\left[-1.95,-1.85\right]$, around the predictions from $N$-body simulation and semi-analytic models $\alpha\sim-1.9$ [Giocoli++08, Springel++08, Fiacconi++16, Benson20, Gannon++25]. We draw $\Sigma_{\rm{sub}}/\rm{kpc^{-2}}$ from a log-uniform sampling distribution in $\log_{10}\Sigma_{\rm{sub}}\sim\left[-2.2,0.2\right]$, which spans a broad range around the theoretical predictions from the Symphony suite of $N$-body simulations $\Sigma_{\rm{sub}}\sim 0.05\ \rm{kpc^{-2}}$ [Nadler++23, Gannon++25] and from the galacticus semi-analytic model, $\Sigma_{\rm{sub}}\sim 0.1\ \rm{kpc^{-2}}$ [Du++25]. Relative to previous analyses [Gilman++20, Keeley++24], we extend the sampling distribution to significantly higher infall mass function amplitudes than these theoretical predictions to test the dependence of the results on the prior, and to demonstrate how imaging data helps break covariance between $\Sigma_{\rm{sub}}$ and $m_{\rm{hm}}$. Finally, for each lens we sample the host halo mass from a log-normal prior $\log_{10}m_{\rm{host}}/M_{\odot}\sim\mathcal{N}\left(13.3,0.3\right)$, based on the typical host halo masses for strong lens systems [Lagattuta++10]. We marginalize over $m_{\rm{host}}$ for each lens before combining likelihoods, meaning the host halo mass is treated as a nuisance parameter in Equation 3.
@@ VI.2 Constraints on WDM
From Figure 19 we see that $\Sigma_{\rm{sub}}$ and $m_{\rm{hm}}$ are correlated. The physical interpretation of this correlation is that adding additional more-massive subhalos to warmer (larger $m_{\rm{hm}}$) dark matter models can mimic the lensing signal of a colder dark matter model with fewer subhalos. Incorporating constraints from imaging data aids in breaking this covariance, strengthening the bound on $m_{\rm{hm}}$ by 0.8 dex, as summarized in the figure caption.
@@
NN-body simulations and semi-analytic models make theoretical predictions for the normalization of the subhalo mass function. However, $N$-body simulations and the semi-analytic model galacticus differ in their predictions by a factor of $\sim 2$. The $N$-body simulations presented by Nadler++23 predict $\Sigma_{\rm{sub}}\sim 0.05\ \rm{kpc^{-2}}$ [Gannon++25], while the semi-analytic model galacticus predicts $\Sigma_{\rm{sub}}\sim 0.1\ \rm{kpc^{-2}}$ [Gannon++25, Du++25]. The priors on subhalo abundance from $N$-body simulations are obtained from the latest generation of subhalo finders [Mansfield++24] applied to the Symphony suite of simulations [Nadler++23]. The galacticus semi-analytic model was calibrated to the caterpillar suite of $N$-body simulations [Griffen++16].
The right panel of Figure 19 shows the posterior distribution with informative priors on $\Sigma_{\rm{sub}}$ based on $N$-body simulations (purple) and semi-analytic model galacticus (green), as indicated in the figure caption. These priors break covariance between $\Sigma_{\rm{sub}}$ and the half-mode mass $m_{\rm{hm}}$. Assuming a prior $\pi\left(\log_{10}\Sigma_{\rm{sub}}/\rm{kpc^{-2}}\right)=\mathcal{N}\left(-1.0,0.3\right)$, as predicted by galacticus, the bound on $m_{\rm{hm}}$ strengthens to $m_{\rm{hm}}<10^{7.4}\mathrm{M}_{\odot}$. This constraint is 0.4 dex stronger than the limit obtained without using lensed arcs, presented in a companion paper, Keeley++25. Assuming $\pi\left(\log_{10}\Sigma_{\rm{sub}}/\rm{kpc^{-2}}\right)=\mathcal{N}\left(-1.3,0.3\right)$, as predicted by $N$-body simulations, we infer $m_{\rm{hm}}<10^{7.2}\mathrm{M}_{\odot}$, an improvement of 0.3 dex relative to omitting imaging data. These upper limits on $m_{\rm{hm}}$ correspond to (see Equation 18) lower bounds on the mass of a thermal relic WDM particle of $7.4\ \rm{keV}$ and $8.4\ \rm{keV}$ for the galacticus and $N$-body priors, respectively. The corresponding one-sided $95\%$ exclusion limits are $m_{\rm{hm}}<10^{6.9}M_{\odot}$ and $m_{\rm{hm}}<10^{6.7}M_{\odot}$, respectively, which correspond to thermal relic particle constraints $m_{\rm{thermal}}>10.4\ \rm{keV}$ and $m_{\rm{thermal}}>11.9\ \rm{keV}$, respectively.

### v2

@@ Abstract
We present a measurement of the free-streaming length of dark matter (DM) and subhalo abundance around 28 quadruple image strong lenses using observations from JWST MIRI presented in Paper III of this series. We improve on previous inferences on DM properties from lensed quasars by simultaneously reconstructing extended lensed arcs with image positions and relative magnifications (flux ratios). Our forward modeling framework generates full populations of subhalos, line-of-sight halos, and globular clusters, uses an accurate model for subhalo tidal evolution, and accounts for free-streaming effects on halo abundance and concentration. Modeling lensed arcs leads to more-precise model-predicted flux ratios, breaking covariance between subhalo abundance and the free-streaming scale parameterized by the half-mode mass $m_{\rm{hm}}$. Assuming subhalo abundance predicted by the semi-analytic model galacticus ($N$-body simulations), we infer (Bayes factor of 10:1) $m_{\rm{hm}}<10^{7.4}\mathrm{M}_{\odot}$ ($m_{\rm{hm}}<10^{7.2}\mathrm{M}_{\odot}$), a 0.4 dex (0.3 dex) improvement relative to omitting lensed arcs. These bounds correspond to lower limits on thermal relic DM particle masses of $7.4$ and $8.4$ keV, respectively. Conversely, assuming DM is cold, we infer a projected mass in subhalos ($10^{6}<m/M_{\odot}<10^{10.7}$) of $1.6_{-1.1}^{+2.4}\times 10^{7}\ \mathrm{M}_{\odot}\ \rm{kpc^{-2}}$ at $95\%$ confidence. This is consistent with galacticus predictions ($0.6\times 10^{7}\mathrm{M}_{\odot}\ \rm{kpc^{-2}}$), but in mild tension with recent $N$-body simulations ($0.3\times 10^{7}\mathrm{M}_{\odot}\ \rm{kpc^{-2}}$). Our results are among the strongest bounds on WDM, and the most precise measurement of subhalo abundance around strong lenses. Further improvements will follow from the large sample of lenses to be discovered by Euclid, Rubin, and Roman.
@@ II.2.2 The flux ratio likelihood
We apply the ABC approach to one system in our sample, J0924, because the flux ratios deviate strongly from the predictions of a smooth lens model. As a result, the flux ratio likelihood heavily downweights the majority of lens models. To generate enough samples to fully explore the parameter space we use the ABC approximation of the exact flux ratio likelihood. With ABC, we can guarantee that we have enough samples to explore the parameter space. We accept the 6,000 samples corresponding to the smallest $S$ out of $\sim 6,000,000$ total lens models. This amounts to an implicit choice of $\epsilon$.
@@ II.3 Imaging data reconstruction
There is not a unique way to choose $\sigma_{\rm{noise}}$. We have experimented with defining this term in a way that connects directly to the uncertainties in the imaging data on a pixel level. For example, we can generate 10,000 background noise realizations (using $\sigma_{ii,(\rm{rms})}$ and $\sigma_{ii,\rm{(Poisson)}}$), render them on top of a best-fit lens model for each system, and set $\sigma_{ii,(\rm{noise})}$ equal to the standard deviations of the resulting distribution of log-likelihoods. However, strategies for defining $\sigma_{(\rm{noise})}$ in a way that connects to the background and Poison noise in the image tends to over-correct, in the sense that we discard more information from the imaging data that is necessary to make the problem computationally tractable. Instead, since the primary motivation for including $\sigma_{(\rm{noise})}$ is related to available CPU hours, $n_{\rm{cpu-hour}}$, we define $\sigma_{(\rm{noise})}$ in such a way that $\sigma_{(\rm{noise})}\rightarrow 0$ (meaning we do not down-weight constraints from lensed arcs) as $n_{\rm{cpu-hour}}\rightarrow\infty$. To compute $w_{\rm{img}}$, we generate $N\sim 200,000$ lens models per system, on average, and set $\sigma_{(\rm{noise})}$ such that $N/N_{\rm{eff}}=2000$. The threshold $N/N_{\rm{eff}}=2000$ represents a compromise between extracting more information from the lensed arcs, and having enough sampling of the parameter space to suppress shot-noise in the $w_{\rm{img}}$ distributions. We offer a complementary interpretation of the $\sigma_{(\rm{noise})}$ parameter: this parameter effectively bootstraps the uncertainties in the imaging data until the top $\sim 1\%$ of lens models have comparable log-likelihoods.
@@ III Datasets
Table 1: Description of the dark matter hyper-parameters introduced in Section IV. The predictions for $\delta_{\rm{LOS}}$, $\alpha$ and $\Sigma_{\rm{sub}}$ remain unchanged in WDM, as these parameters determine the form of the subhalo mass function anchored at high masses.
@@ IV.1.2 Subhalos
Our strategy for modeling subhalo density profiles uses the tidal evolution model presented by Du++25. This model is designed specifically for the analysis of strong gravitational lens systems, and predicts the density profile of subhalos that appear in projection inside $\sim 3\theta_{\rm{E}}$, or approximately $\sim 30\ \rm{kpc}$ for a typical system, based on the subhalo properties at infall, given the infall subhalo mass function.
We begin by defining the differential subhalo mass function per unit area at infall
@@
Equation 15 specifies the number density of halos with a mass $m$ at infall that eventually appear within a $30\ \rm{kpc}$ aperture of the host halo center (the region relevant for strong lensing). After generating this infall population, we use the tidal stripping framework presented by Du++25 to predict each subhalo’s density profile. The calculation begins with assigning infall redshifts to each subhalo using the probability distribution $p\left(z_{\rm{infall}}|m,m_{\rm{host}}\right)$, where the infall redshift depends on $m$ and the host-halo mass $m_{\rm{host}}$. From $z_{\rm{infall}}$ and $m$ we evaluate the concentration of the subhalo at infall $c=c\left(m,z_{\rm{infall}}\right)$. Next, given the host halo concentration $c_{\rm{host}}=c_{\rm{host}}\left(m_{\rm{host}},z_{\rm{d}}\right)$, which we calculate using the DiemerJoyce19 concentration–mass relation, we sample the bound mass fraction, $f_{\rm{bound}}\equiv m_{\rm{bound}}/m$, from the probability distribution $p\left(f_{\rm{bound}}|c,z_{\rm{infall}},c_{\rm{host}}\right)$. Du++25 derived empirical representations of this distribution by evolving subhalos in a growing host potential with galacticus. As discussed by Du++25, a given set of subhalo infall time, concentration, and host halo concentration specifies a distribution of $f_{\rm{bound}}$ due to differences in subhalo orbits and mass loss histories. Given $f_{\rm{bound}}$, we use tidal tracks [e.g. Errani++21, Du++24] to calculate the truncation radius $r_{\mathrm{t}}$, and a factor $f_{\rm{t}}$ that changes the normalization of the NFW profile $\rho_{\mathrm{s}}\rightarrow f_{\rm{t}}\rho_{\mathrm{s}}$. This model predicts that, at the redshift of the lens, most subhalos have lost over $90\%$ of the mass they had at infall. The model implemented in $\tt{pyHalo}$ shows excellent agreement with cosmological simulations of subhalo evolution performed with galacticus for $\Sigma_{\rm{sub}}\approx 0.1\ \rm{kpc^{-2}}$ [Du++25].
@@ IV.2 Extension to warm dark matter
The values of $k_{\rm{hm}}$, $\beta$, and $\gamma$ depend on the mass and formation mechanism of the dark matter particle. For example, a sterile neutrino of the same mass can be warm or cold, depending on the production mechanism [Kusenko:2009up, Abazajian:2019ejt, Zelko++22]. To enable comparison with existing WDM constraints, for this initial analysis of our full JWST dataset we assume the same thermal relic WDM model considered in many previous works. For a thermal relic WDM particle [^15] For thermal relic WDM we have $\beta\sim 2$ and can take $\gamma=5$ in Equation 17 [Stucker++22]., we can relate the half-mode mass to the particle mass
@@
We note that Nadler++25cozmic present a revised version of this formula $m_{\rm{hm}}=4.3\times 10^{8}\left(\frac{m_{\rm{therm}}}{3\rm{keV}}\right)^{-3.564}\mathrm{M}_{\odot}$. We will use Equation 18 when comparing with previous work, but we will also quote results using the modified version in Section VI.
@@
IV.3 Sampling distributions and priors for dark matter hyper-parameters
We use a uniform sampling distribution on $\log_{10}m_{\rm{hm}}/\mathrm{M}_{\odot}$ in the range $\left[4,10\right]$, and quote our constraints on this parameter with the same Bayesian prior. While CDM encompasses models with halos less massive than $10^{4}M_{\odot}$, and therefore $m_{\rm{hm}}=10^{4}\mathrm{M}_{\odot}$ is not formally consistent with the theory, for $m_{\rm{hm}}=10^{4}\mathrm{M}_{\odot}$ the abundance and density profiles of halos deviate by less than $10\%$ at $10^{6}\mathrm{M}_{\odot}$, the minimum halo mass rendered in our simulations. Therefore, we can identify regions of parameter space with $m_{\rm{hm}}<\mathcal{O}\left(10^{5}\mathrm{M}_{\odot}\right)$ as effectively indistinguishable from CDM.
We draw $\delta_{\rm{LOS}}$ (Equation 13) from a sampling distribution that is uniform in $\left[0.9,1.1\right]$, allowing for $10\%$ difference in the amplitude of the line-of-sight halo mass function, relative to the Sheth-Tormen model. We draw $\alpha$ from a uniform sampling distribution in $\left[-1.95,-1.85\right]$, around the predictions from $N$-body simulation and semi-analytic models $\alpha\sim-1.9$ [Giocoli++08, Springel++08, Fiacconi++16, Benson20, Gannon++25].
We draw $\Sigma_{\rm{sub}}/\rm{kpc^{-2}}$ from a log-uniform sampling distribution in $\log_{10}\Sigma_{\rm{sub}}\sim\left[-2.2,0.2\right]$, which spans a broad range around the theoretical predictions from the Symphony suite of $N$-body simulations. The value of $\Sigma_{\rm{sub}}$ that reproduces the galacticus bound mass function, given our tidal evolution model, and an additional suppression of $15\%$ due to the presence of the central elliptical galaxy [Gannon++25], is $\Sigma_{\rm{sub}}=0.1\ \rm{kpc^{-2}}$ [Du++25]. The bound mass function predicted by Symphony is lower by a factor of 2 relative to the galacticus bound mass function [Gannon++25]. This implies $\Sigma_{\rm{sub}}\sim 0.05\ \rm{kpc^{-2}}$ as the prediction from recent Symphony N-body simulations [Nadler++23]. In Section VI, we will present constraints on $m_{\rm{hm}}$ that use priors centered on these values with a factor of 2 uncertainty. We note that the analysis by Gannon++25 used the public dataset from the Symphony simulations, which makes use of the latest generation of halo finders [Mansfield++24]. The galacticus semi-analytic model was calibrated to the caterpillar suite of $N$-body simulations [Griffen++16].
Relative to previous analyses [Gilman++20, Keeley++24], we extend the sampling distribution on $\Sigma_{\rm{sub}}$ to significantly higher infall mass function amplitudes than these theoretical predictions to test the dependence of the results on the prior, and to demonstrate how imaging data helps break covariance between $\Sigma_{\rm{sub}}$ and $m_{\rm{hm}}$. Finally, for each lens we sample the host halo mass from a log-normal prior $\log_{10}m_{\rm{host}}/M_{\odot}\sim\mathcal{N}\left(13.3,0.3\right)$, based on the typical host halo masses for strong lens systems [Lagattuta++10]. We marginalize over $m_{\rm{host}}$ for each lens before combining likelihoods, meaning the host halo mass is treated as a nuisance parameter in Equation 3.
@@ VI.2 Constraints on WDM
From Figure 19 we see that $\Sigma_{\rm{sub}}$ and $m_{\rm{hm}}$ are correlated. The physical interpretation of this correlation is that adding additional more-massive subhalos to warmer (larger $m_{\rm{hm}}$) dark matter models can mimic the lensing signal of a colder dark matter model with fewer subhalos. Incorporating constraints from imaging data aids in breaking this covariance, strengthening the bound on $m_{\rm{hm}}$ by 0.7 dex, as summarized in the figure caption.
@@
NN-body simulations and semi-analytic models make theoretical predictions for the normalization of the subhalo mass function. However, $N$-body simulations and the semi-analytic model galacticus differ in their predictions by a factor of $\sim 2$. The $N$-body simulations presented by Nadler++23 predict $\Sigma_{\rm{sub}}\sim 0.05\ \rm{kpc^{-2}}$ [Gannon++25], while the semi-analytic model galacticus predicts $\Sigma_{\rm{sub}}\sim 0.1\ \rm{kpc^{-2}}$ [Gannon++25, Du++25] (see discussion in Section IV.3).
The right panel of Figure 19 shows the posterior distribution with informative priors on $\Sigma_{\rm{sub}}$ based on $N$-body simulations (purple) and semi-analytic model galacticus (green), as indicated in the figure caption. These priors break covariance between $\Sigma_{\rm{sub}}$ and the half-mode mass $m_{\rm{hm}}$. Assuming a prior $\pi\left(\log_{10}\Sigma_{\rm{sub}}/\rm{kpc^{-2}}\right)=\mathcal{N}\left(-1.0,0.3\right)$, as predicted by galacticus, the bound on $m_{\rm{hm}}$ strengthens to $m_{\rm{hm}}<10^{7.4}\mathrm{M}_{\odot}$. This constraint is 0.4 dex stronger than the limit obtained without using lensed arcs, presented in a companion paper, Keeley++25. Assuming $\pi\left(\log_{10}\Sigma_{\rm{sub}}/\rm{kpc^{-2}}\right)=\mathcal{N}\left(-1.3,0.3\right)$, as predicted by $N$-body simulations, we infer $m_{\rm{hm}}<10^{7.2}\mathrm{M}_{\odot}$, an improvement of 0.3 dex relative to omitting imaging data. These upper limits on $m_{\rm{hm}}$ correspond to (see Equation 18) lower bounds on the mass of a thermal relic WDM particle of $7.4\ \rm{keV}$ and $8.4\ \rm{keV}$ for the galacticus and $N$-body priors, respectively. The corresponding one-sided $95\%$ exclusion limits are $m_{\rm{hm}}<10^{6.9}M_{\odot}$ and $m_{\rm{hm}}<10^{6.7}M_{\odot}$, respectively, which correspond to thermal relic particle constraints $m_{\rm{thermal}}>10.4\ \rm{keV}$ and $m_{\rm{thermal}}>11.9\ \rm{keV}$, respectively. As discussed in Section IV.2 (immediately following Equation 18) a slight modification to the conversion between $m_{\rm{hm}}$ and $m_{\rm{therm}}$ presented by Nadler++25cozmic gives bounds of $6.7$ keV and $7.6$ keV for the two Bayes factor limits.